OpenStudy (anonymous):

In the following system of equations, what does x equal? -8x – 10y = -36 -8x – 6y = -28

jhonyy9 (jhonyy9):

-8x -10y =-36 multiplie by -1 and add to second 8x +10y =36 -8x -6y =-28 ---------------- 0 4y=8 --- y=2 -8x-10y=-36 -8x-10*2=-36 -8x-20=-36 -8x=-16 x=2 x=2 y=2

jhonyy9 (jhonyy9):

ok ?